Night begins with Rex telluric and science (1:06) (1:30)

Wiphu starts (3:50) and completes his program (5:30) seeing variable from 1.5-3"

Back to Rex (5:45)

seeing still variable 1.5-2" (7:35)

Altering McGreer's scripts to use 2" slit for poor seeing.

Altering McGreer's telluric script to use brighter guide star with rotation 240. (10:00) (j0353)

Altering McGreer's blindoffset script to new position angle, new guide star - otherwise guidestar fails after large blind offset

started mcgreer's science (~11:35)

wavefront sensing diverging --> cleared forces on mirror - this was done during frame 0176 (12:26)

collimation still deteriorating, frames 176 onward probably not usable, the few previous frames may or may not be useful --> aborted script

Dark script on 20101102UT failed and no files were saved (flush detector command in script causes script to fail - must comment out or delete this in dark scripts!) - Consequently darks for 20101102UT integrations and 20101103UT integrations were all taken tonight.

-- EmilyMcLinden - 03 Nov 2010
Topic revision: r4 - 03 Nov 2010, EmilyMcLinden
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback